# Toward Learning POMDPs Beyond Full-Rank Actions and State Observability

### v1

@@ B.2 Runtime Complexity in Floating-Point Operations
Interestingly, our calculation suggests a runtime that is polynomial when the observability length $n^{\mathit{obs}}$ is known to be bounded by a constant. Further investigation to prove fixed-parameter tractability under a computational learning framework (e.g. PAC-learning) would be an interesting direction of future work.
Appendix C Additional Experimental Details
@@ C.1 Algorithm Parameter Selection
For Hankel estimation, of practical concern is the selection of the size Hankel matrix to estimate, or the sequences to include as row and column indices. Like many other approaches (hsuSpectralAlgorithm2012; balleSpectralLearning2014), we that we use every possible action-observation sequence up to a certain length. We expose this length as an algorithm parameter. While smaller lengths will result in faster convergence of matrix entry estimates, selecting a length that is too short may result in a Hankel matrix whose approximate rank is strictly less than the number of states of the system. Our chosen lengths are included in Table 1. Automatically determining the proper Hankel size is an open question since the development of spectral approaches for PSRs (wolfeLearningPredictive2005; bootsClosingLearningplanning2011; balleSpectralLearning2014), and remains an interesting question for future work.
@@
When recovering the observation distributions and partition-level transitions, we must specify thresholds to determine transition matrix full-rankness and a threshold that determines when observation distributions are similar (Section B.1). Inverting a near-singular matrix is highly undesirable when computing the matrices to joint-diagonalize in Eq. 17, so we specify conservatively high threshold on the smallest singular value on the smallest singular value $\sigma_{min}$ of the transition matrix. As discussed in Appendix B.1, in practice, a threshold $\tau_{obs}$ is not required to compute a random block-diagonal rotation matrix $R$. However, $\tau_{obs}$ is still required to compute partition-level transition likelihoods errors reported in Figures 3 and 4. We set a conservativitely high threshold $\tau_{obs}$ to merge observation distributions aggressively when plotting those figures.
@@ C.3 Planning Performance of Different Sampling and Distribution Rounding Strategies
There are many ways to sample action-observation trajectories when deriving a UCT-based search algorithm for planning on POMDPs. As discussed by silver_montecarlo_2010, there are largely two approaches, which differ in how the latent state treated as the search propogates down a branch of the search tree.
@@
    Upon choosing an action, propogate the full belief state using a process update (e.g. multiplying by $\tilde{T}^{a}$ of Theorem 1). Compute the mixture observation distribution weighted by that belief state, from which we can sample the emitted observation.
@@
    Upon choosing an action, sample a single latent state (or observability partition, in the context of Theorem 1). Look up the observation distribution associated with the action and sampled state, and then sample the observation.
silver_montecarlo_2010 prove that the observation distribution under these two sampling strategies are equivalent, so a UCT-based search will perform the same using either approach. They also argue the latter is more computationally efficient for systems with a large number of states.
@@
where $\mathbb{I}_{S}$ is a vector with entries of value one for indices in partition $S$ and zero otherwise, $a$ is the selected action by the search algorithm, and $\otimes$ the element-wise product. The conditional observation distribution is then found by computing products $\tilde{b}_{S}\tilde{T}^{a}O^{ao}$ for all observations $o\in\mathcal{O}$, and projecting the distribution to deal with approximation error as handlied in the first appraoch. Verifying the correctness of this calculation is a straightforward extension of the proof of Theorem 1.
@@ C.4 Slower Convergence of Transition Likelihoods
A common approach to convergence analysis of tensor decomposition methods would first argue the convergences of the SVD of our Hankel matrix and then argue convergence of the eigendecompositions of the matrix discussed in 18 (moitraAlgorithmicAspects2018). PSRs only depend on the convergence of the SVD, while our learned POMDPs depends on the convergence of both the SVD and eigendecomposition. In our reward-specification experiments (Fig. 4), accurate transitions are required to correctly assign reward to the desired goal state. The slow convergence of performance of the planner in the directional hallway environment suggests that more data is required to obtain accurate likelihood estimates of transition and observation matrices.
@@ C.5 Experimental Domains
Here, we document any environment we have modified, or any novel environments we introduced in this work. We have used the original Tiger domain as described by kaelbling_planning_1998, which we omit from our discussion below.
@@ C.5.1 Sense-Float-Reset
As discussed in Sec. 4, the transition dynamics and observation emissions of Sense-Float-Reset are the same as those of Float-Reset introduced by littmanPredictiveRepresentations2001, but augment the system with a passive sensing action.
@@ C.5.2 T-Maze
We present a version of T-Maze similar to the one described by allenMitigatingPartial2024. Since we allow actions to determine observation emissions (allenMitigatingPartial2024 determine observations by states), our T-Maze POMDP has fewer states than their version. This environment is more easily explained pictorally than explicit probability expressions. See Fig. 6 for a depiction of transition dynamics and observation emissions. For the truncated T-Maze used for experiments in Section 5 Figure 3, the number of room states was set to $k=1$.
@@ C.5.3 Noisy Hallways
Figure 7: Noisy-hallway dynamics shared between directional hallway and noisy hallway domains. Edges are labeled with transition probabilities, and nonzero rewards are emitted deterministically from annotated states. Self-loop edges with probability less than 11 are omitted. For the reset action, the agent samples the next state uniformly from the available states.
@@
There is no reward function is given, since both of these experiments are used in the reward-specification experiments discussed in Sec. 5, Fig. 4. For the directional environment, the rewarded tuples are $(\texttt{left},\texttt{end-left})$ and $(\texttt{right},\texttt{end-right})$. For the noisy environment, the rewarded tuples are $(\texttt{left},\texttt{end-left})$, $(\texttt{right},\texttt{end-right})$, $(\texttt{left},\texttt{end-right})$, and $(\texttt{right},\texttt{end-left})$.
@@ C.6 Example Output of Algorithm
Here, we include an example of the learned model after estimates of transition and observation matrices have nearly converged. In Tiger, where each state has a unique observation distribution, the learned model, as illustrated in Fig. 8, shows close agreement between the learned and ground-truth transition and observation matrices. These results confirm that by estimating the similarity transform, we can recover the true observation and transition models.
Appendix D Use of Large Language Models
Our use of large language models is solely limited to the completion of routine coding tasks in Python. These routine tasks consist of experiment launching, data retrieval infrastructure, and plotting code.

### v2

@@ B.2 Runtime Complexity in Floating-Point Operations
Interestingly, our calculation suggests a runtime that is polynomial when the observability length $n^{\mathit{obs}}$ scales favorably when as the number of states of a particular class of POMDPs increases, which aligns on prior work on Hidden Markov Models [huangMinimalRealization2016]. Further investigation of computational tractability learning framework (e.g. PAC-learning) for the multi-action case would be an interesting direction of future work.
C Appendix: Additional Experimental Details
@@ C.1 Algorithm Parameter Selection
For Hankel estimation, of practical concern is the selection of the size Hankel matrix to estimate, or the sequences to include as row and column indices. Like many other approaches [hsuSpectralAlgorithm2012, balleSpectralLearning2014], we that we use every possible action-observation sequence up to a certain length. We expose this length as an algorithm parameter. While smaller lengths will result in faster convergence of matrix entry estimates, selecting a length that is too short may result in a Hankel matrix whose approximate rank is strictly less than the number of states of the system. Our chosen lengths are included in Table 1. Automatically determining the proper Hankel size is an open question since the development of spectral approaches for PSRs [wolfeLearningPredictive2005, bootsClosingLearningplanning2011, balleSpectralLearning2014], and remains an interesting question for future work.
@@
When recovering the observation distributions and partition-level transitions, we must specify thresholds to determine transition matrix full-rankness and a threshold that determines when observation distributions are similar (Section B.1). Inverting a near-singular matrix is highly undesirable when computing the matrices to joint-diagonalize in Eq. 15, so we specify conservatively high threshold on the smallest singular value on the smallest singular value $\sigma_{min}$ of the transition matrix. As discussed in Appendix B.1, in practice, a threshold $\tau_{obs}$ is not required to compute a random block-diagonal rotation matrix $R$. However, $\tau_{obs}$ is still required to compute partition-level transition likelihoods errors reported in Figures 4 and 5. We set a conservativitely high threshold $\tau_{obs}$ to merge observation distributions aggressively when plotting those figures.
@@ C.3 Planning Performance of Different Sampling and Distribution Rounding Strategies
There are many ways to sample action-observation trajectories when deriving a UCT-based search algorithm for planning on POMDPs. As discussed by [silver_montecarlo_2010], there are largely two approaches, which differ in how the latent state treated as the search propogates down a branch of the search tree.
@@
    Upon choosing an action, propogate the full belief state using a process update (e.g. multiplying by $\tilde{T}^{a}$ of Theorem 4.1). Compute the mixture observation distribution weighted by that belief state, from which we can sample the emitted observation.
@@
    Upon choosing an action, sample a single latent state (or observability partition, in the context of Theorem 4.1). Look up the observation distribution associated with the action and sampled state, and then sample the observation.
[silver_montecarlo_2010, Lemma 2] prove that the observation distribution under these two sampling strategies are equivalent, so a UCT-based search will perform the same using either approach. They also argue the latter is more computationally efficient for systems with a large number of states.
@@
where $\mathbb{I}_{S}$ is a vector with entries of value one for indices in partition $S$ and zero otherwise, $a$ is the selected action by the search algorithm, and $\otimes$ the element-wise product. The conditional observation distribution is then found by computing products $\tilde{b}_{S}\tilde{T}^{a}O^{ao}$ for all observations $o\in\mathcal{O}$, and projecting the distribution to deal with approximation error as handlied in the first appraoch. Verifying the correctness of this calculation is a straightforward extension of the proof of Theorem 4.1.
@@ C.4 Slower Convergence of Transition Likelihoods
A common approach to convergence analysis of tensor decomposition methods would first argue the convergences of the SVD of our Hankel matrix and then argue convergence of the eigendecompositions of the matrix discussed in 16 [moitraAlgorithmicAspects2018]. PSRs only depend on the convergence of the SVD, while our learned POMDPs depends on the convergence of both the SVD and eigendecomposition. In our reward-specification experiments (Fig. 5), accurate transitions are required to correctly assign reward to the desired goal state. The slow convergence of performance of the planner in the directional hallway environment suggests that more data is required to obtain accurate likelihood estimates of transition and diagonal observation matrices.
@@ C.5 Experimental Domains
Here, we document any environment we have modified, or any novel environments we introduced in this work. We have used the original Tiger domain as described by [kaelbling_planning_1998], which we omit from our discussion below.
@@ C.5.1 Sense-Float-Reset
As discussed in Sec. 4, the transition dynamics and observation emissions of Sense-Float-Reset are the same as those of Float-Reset introduced by [littmanPredictiveRepresentations2001], but augment the system with a passive sensing action.
@@ C.5.2 T-Maze
We present a version of T-Maze similar to the one described by [allenMitigatingPartial2024]. Since we allow actions to determine observation emissions ([allenMitigatingPartial2024] determine observations by states), our T-Maze POMDP has fewer states than their version. This environment is more easily explained pictorally than explicit probability expressions. See Fig. 7 for a depiction of transition dynamics and observation emissions. For the truncated T-Maze used for experiments in Section 5 Figure 4, the number of room states was set to $k=1$.
@@ C.5.3 Noisy Hallways
There is no reward function is given, since both of these experiments are used in the reward-specification experiments discussed in Sec. 5, Fig. 5. For the directional environment, the tuples that are assigned rewards are $(\texttt{left},\texttt{end-left})$ and $(\texttt{right},\texttt{end-right})$. For the noisy environment, the tuples assigned rewards are the elements of the set $\{\texttt{left},\texttt{right}\}\times\{\texttt{end-left},\texttt{end-right}\}$.
@@ C.6 Example Output of Algorithm
Here, we include an example of the learned model after estimates of transition and diagonal observation matrices have nearly converged. In Tiger, where each state has a unique observation distribution, the learned model, as illustrated in Fig. 8, shows close agreement between the learned and ground-truth transition and diagonal observation matrices. These results confirm that by estimating the similarity transform, we can recover the true observation and transition models.